OpenStudy (ksaimouli):

solve for x using the quadratic formula x2+5x-1=0

OpenStudy (anonymous):

now this one you need the formula for

OpenStudy (anonymous):

use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=1,b=5,c=-1\] you get \[x=\frac{-5\pm\sqrt{5^2-4\times 1\times -1}}{2\times1}\] \[\frac{-5\pm\sqrt{25+4}}{2}\] \[\frac{-5\pm\sqrt{29}}{2}\]

OpenStudy (ksaimouli):

after this

OpenStudy (anonymous):

He gave you the answers. There are two. Compute -5+root(29)/2 and -5-root(29)/2

OpenStudy (anonymous):

more than this there is nothing